If $20,000 is invested in a savings account offering 3.5% per year, compounded continuously, how fast is the balance growing after 4 years? (Round your answer to the nearest cent.)
$ 1 per yr

Answer by rfer(16322)

A=20000(2.7183)^4*0.035
A=20000(2.7183)^0.14
A=20000*1.15
A=23,000.00